OpenStudy (itsbribro):

Please help i dont understand!!! Which is an equation in point-slope form for the given point and slope? Point: (8,-4) ; Slope: -5/6

OpenStudy (anonymous):

y=mx+c is the point slope form. You have c=-5/6 Hence re write the equation as y=mx-5/6. Next put the values of x=8 and y=-4 and compute m . since you then get both m and c, just write the final answer in terms of x and y